# Limits to extreme event forecasting in chaotic systems

### v1

@@ Abstract
Predicting extreme events in chaotic systems, characterized by rare but intensely fluctuating properties, is of great importance due to their impact on the performance and reliability of a wide range of systems. Some examples include weather forecasting, traffic management, power grid operations, and financial market analysis, to name a few. Methods of increasing sophistication have been developed to forecast events in these systems. However, the boundaries that define the maximum accuracy of forecasting tools are still largely unexplored from a theoretical standpoint. Here, we address the question: What is the minimum possible error in the prediction of extreme events in complex, chaotic systems? We derive lower bounds for the minimum probability of error in extreme event forecasting using the information-theoretic Fano’s inequality. The limits obtained are universal, in that they hold regardless of the modeling approach: from traditional linear regressions to sophisticated neural network models. The approach also allows us to assess whether reduced-order models are operating near their theoretical maximum performance or if further improvements are theoretically possible.
@@ 1 Introduction
Extreme events, characterized by rare but intensely fluctuating properties, are ubiquitous in both engineering system and natural phenomena Albeverio2006 . For instance, turbulent gusts over an aircraft can result in bumpy flights Etkin1981 , severe weather can disrupt communication systems Rahmstorf2011 , rare but large cascades in electrical power grids may lead to failures Schaefer2018 , and sudden increases in traffic flow can trigger network paralysis Kumar2020 . In these scenarios, the real-time prediction of extreme events is crucial for enabling proactive measures to avert potential issues Farazmand2019 ; Sapsis2021 . By accurately forecasting the extreme states of dynamical systems, we can mitigate adverse effects, reduce downtime, and prevent failures. In this study, we present the lower bound on the accuracy of extreme event prediction using the framework of information theory. The bound obtained is a fundamental limit –independent of the modeling approach– that arises from the finite amount of information the observed state contains about the extreme event.
A variety of methods have been employed to predict extreme events in time series of chaotic dynamical systems. Some of the approaches that have proven effective include nonlinear dynamics estimation based on the Koopman operator theory Otto2021 and Takens embedding theorem Asch2022 , along with machine learning techniques, such as support vector machines Mukherjee1997 , singular spectrum analysis and the maximum entropy method Ghil2002 . Advanced deep learning methods, including auto-encoders Racah2017 , long short-term memory networks Vlachas2018 , and reservoir computing Pammi2023 have also been instrumental to devise forecasting models for chaotic systems with high-dimensional attractors.
@@
Improvements in the prediction of extreme events can be achieved either by enhancing models to better represent the physics, gaining access to more observables, or reducing uncertainty in the initial conditions. Eliminating modeling errors is theoretically possible, given the knowledge of a set of governing equations that reflect the underlying dynamics of the observed system. However, accessing variables beyond what is currently observable may be limited by experimental/computational constraints. Additionally, no feasible approach can completely eliminate prediction errors caused by uncertainty in the initial conditions. Even with highly precise measurements, minor errors in the initial state eventually amplify, compromising the accuracy of the forecast for long times (Lorenz1963, ). Here, our focus is not on developing superior models for extreme event prediction. Instead, we pose the fundamental question: what is the theoretically maximum achievable accuracy in extreme event prediction regardless of the modeling approach and source of error?
2 Methods
Consider the quantity of interest $Q(t)$ from a chaotic system, where $t$ is time. The extreme event indicator $E(t)$ is defined as
$E(t)=\begin{cases}1&\text{ if }Q(t)\geqslant\eta,\\ 0&\text{otherwise},\end{cases}$ (1)
where $\eta$ is the threshold for extreme event detection. The vector of observable variables is $\boldsymbol{\hat{Q}}(t)=[\hat{Q}_{1}(t),\hat{Q}_{2}(t),...]$, which contains the accessible information about the system. One could forecast the extreme event $E(t)$ at time $t$ using the finitely precise observable in the past, $\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{Q}}(t-\delta t)\pm\delta\boldsymbol{\hat{Q}}$, as the input to the model $\hat{f}$ such that $\hat{E}(t)=\hat{f}\left(\boldsymbol{\hat{Q}}^{-}\right)$, where $\hat{E}$ is the modeled extreme event indicator (which might differ from $E$), $\delta t>0$ is the time horizon for extreme event forecasting, and $\delta\boldsymbol{\hat{Q}}$ is the uncertainty in the state $\boldsymbol{\hat{Q}}$. The performance of the model can be evaluated using the probability of error $P_{e}(\boldsymbol{\hat{Q}}^{-},\hat{f})=\text{Probability}(\hat{E}\neq E)=P(\hat{f}(\boldsymbol{\hat{Q}}^{-})\neq E)$.
@@ 2 Methods
Mispredicted extreme events can manifest as either false positives, $P(\hat{E}=1,E=0)$, or false negatives, $P(\hat{E}=0,E=1)$. However, these two types of errors can bear significantly different consequences. For instance, incorrectly predicting a hurricane (false positive) might be inconvenient, but failing to predict one (false negative) can be catastrophic. To accurately reflect the distinct impact of false positive and negative, we introduce the cost-sensitive probability of error:
@@
where $c^{+}$ and $c^{-}$ are the false positive and negative weighting factors, respectively.
The goal is to estimate the minimum cost-sensitive probability of error given the observable state $\boldsymbol{\hat{Q}}^{-}$ over all possible models $\hat{f}$,
@@
The factors $c^{-}$ and $c^{+}$ are scaled such that $21/c^{+}+1/c^{-}=$ 2 to ensure that the range of $P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-})$ is [0, 0.5]. The key idea is that the prediction of extreme events can be intuitively understood as an information transmission process, where information from the current observable state is conveyed to predict the future state (LozanoDuran2022, ). If the forecast is treated as a noisy channel, then the Fano’s inequality (Fano1961, ) provides the foundations for deriving a bound on the minimum probability of error in the transmission of discrete messages. First, we measure the uncertainty in the extreme event indicator $E$ after knowing the present observable $\boldsymbol{\hat{Q}}^{-}$ using the cost-sensitive, conditional Shannon entropy (Shannon1948, ; Zhao2013, )
$H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ $=\sum_{\boldsymbol{\hat{q}}^{-}}h^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),$ (4)
where $\boldsymbol{\hat{q}}^{-}$ is a particular state for $\boldsymbol{\hat{Q}}^{-}$, $P(E\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ is the conditional probability of $E$ on $\boldsymbol{\hat{Q}}^{-}$, and $h^{c}(p)=h\left(\min\{c^{-}p,c^{+}(1-p)\}\right)$ is the cost-sensitive binary entropy function, where $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$. For $c^{+}=c^{-}=1$, $H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ is equal to the classic Shannon conditional entropy. Equation (4) quantifies the additional information required to determine the outcome of $E$ given the information in $\boldsymbol{\hat{Q}}^{-}$ accounting for the weighting factors $c^{-}$ and $c^{+}$. With that, the minimum cost-sensitive probability of error attainable by any model is [see Appendix A]
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=h^{-1}\left(H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right).$ (5)
The lower bound given by Eq. (5) is the consequence of the unavoidable uncertainty intrinsic to chaotic systems. This uncertainty arises from unobserved variables and errors in the initial condition values, which transcend the predictive capabilities of any model. It is worth noting that while the bound in Eq. (5) is utilized here to establish limits for predicting extreme events, it is generally applicable to the prediction of any binary events.

### v2

@@ Abstract
Predicting extreme events in chaotic systems, characterized by rare but intensely fluctuating properties, is of great importance due to their impact on the performance and reliability of a wide range of systems. Some examples include weather forecasting, traffic management, power grid operations, and financial market analysis, to name a few. Methods of increasing sophistication have been developed to forecast events in these systems. However, the boundaries that define the maximum accuracy of forecasting tools are still largely unexplored from a theoretical standpoint. Here, we address the question: What is the minimum possible error in the prediction of extreme events in complex, chaotic systems? We derive the minimum probability of error in extreme event forecasting along with its information-theoretic lower and upper bounds. These bounds are universal for a given problem, in that they hold regardless of the modeling approach for extreme event prediction: from traditional linear regressions to sophisticated neural network models. The limits in predictability are obtained from the cost-sensitive Fano’s and Hellman’s inequalities using the Rényi entropy. The results are also connected to Takens’ embedding theorem using the information can’t hurt inequality. Finally, the probability of error for a forecasting model is decomposed into three sources: uncertainty in the initial conditions, hidden variables, and suboptimal modeling assumptions. The latter allows us to assess whether prediction models are operating near their maximum theoretical performance or if further improvements are possible. The bounds are applied to the prediction of extreme events in the Rössler system and the Kolmogorov flow.
@@ 1 Introduction
Extreme events, characterized by rare but intensely fluctuating properties, are ubiquitous in both engineering system and natural phenomena Albeverio2006 . For instance, turbulent gusts over an aircraft can result in bumpy flights Etkin1981 , severe weather can disrupt communication systems Rahmstorf2011 , rare but large cascades in electrical power grids may lead to failures Schaefer2018 , extreme ocean temperature oscillations could impact agriculture and ecosystems ray2020understanding , rare but significant fluctuations in brain network could cause seizures ray2022extreme , and sudden increases in traffic flow can trigger network paralysis Kumar2020 . In these scenarios, the real-time prediction of extreme events is crucial for enabling proactive measures to avert potential issues Farazmand2019 ; Sapsis2021 . By accurately forecasting the extreme states of dynamical systems, we can mitigate adverse effects, reduce downtime, and prevent failures. In this study, we investigate the limits of predictability in extreme event detection using the framework of information theory. The limit obtained is a fundamental property – independent of the modeling approach – that arises from the finite amount of information the observed state contains about the extreme event.
A variety of methods have been employed to predict extreme events in time series of chaotic dynamical systems. Some of the approaches that have proven effective include nonlinear dynamics estimation based on the Koopman operator theory Otto2021 and Takens embedding theorem Asch2022 , along with machine learning techniques, such as support vector machines Mukherjee1997 , singular spectrum analysis and the maximum entropy method Ghil2002 . Advanced deep learning methods, including auto-encoders Racah2017 , long short-term memory networks Vlachas2018 , and reservoir computing Pammi2023 have also been instrumental to devise forecasting models for chaotic systems with high-dimensional attractors. A discussion on the role of information in the context of model prediction and control for chaotic dynamical systems can be found in Ref. LozanoDuran2022 .
@@
Improvements in the prediction of extreme events can be achieved either by enhancing models to better represent the physics, gaining access to more observables, or reducing uncertainty in the initial conditions. Eliminating modeling errors is theoretically possible, given the knowledge of a set of governing equations that reflect the underlying dynamics of the observed system. However, accessing variables beyond what is currently observable may be limited by experimental or computational constraints. Additionally, no feasible approach can completely eliminate prediction errors caused by uncertainty in the initial conditions. Even with highly precise measurements, minor errors in the initial state eventually amplify due to chaos, compromising the accuracy of the forecast for long times (Lorenz1963, ). Here, our focus is not on developing superior models for extreme event prediction. Instead, we pose the fundamental question: what is the theoretically maximum achievable accuracy in extreme event prediction regardless of the modeling approach and source of error?
2 Formulation
2.1 Modeled extreme event indicator
Consider a chaotic dynamical system completely determined by $N$ time-dependent variables given by the vector $\boldsymbol{Q}(t)=[Q_{1}(t),Q_{2}(t),\dots,Q_{N}(t)]$, where $t$ is the time. We are interested in the extreme values of the variable $Q_{E}(t)$, which is a function of $\boldsymbol{Q}(t)$. The extreme event indicator $E(t)$ is defined as
$E(t)=\begin{cases}1&\text{ if }Q_{E}(t)\geqslant\eta,\\ 0&\text{otherwise},\end{cases}$ (1)
where $\eta$ is the threshold for extreme event detection. The specific value of the threshold $\eta$ is dependent on the problem and could be selected based on the definition of extreme event for each particular application. The vector of observable variables is defined as $\boldsymbol{\hat{Q}}(t)=[\hat{Q}_{1}(t),\hat{Q}_{1}(t),\dots,\hat{Q}_{M}(t)]$, which contains the accessible information about the system (i.e., the variables that can be measured or are assumed to be known). The components of $\boldsymbol{\hat{Q}}$ correspond to individual components of $\boldsymbol{Q}$ or functions of them. In general, $M\leq N$ and the number of observed variables $M$ is equal or smaller than the number of degrees of freedom of the system $N$.
We aim to build a predictive model for $E$. To that end, we define the limited-precision observable containing information from the present time and $p\geq 0$ times in the past:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\hat{Q}}^{-}=[\boldsymbol{\hat{Q}}(t),\boldsymbol{\hat{Q}}(t-\delta t_{1}),\dots,\boldsymbol{\hat{Q}}(t-\delta t_{p})]\pm\delta\boldsymbol{\hat{Q}}^{-},}$ (2)
where $\delta t_{i}>0,i=0,1,\dots,p$ are the time lags used for prediction, and $\delta\boldsymbol{\hat{Q}}^{-}$ is the uncertainty in the observations. The latter may arise experimentally from inaccuracies in measurement tools, numerically from round-off errors in $\boldsymbol{\hat{Q}}$ or its discretization, and generally, from any uncertainties in the value of $\boldsymbol{\hat{Q}}^{-}$. One could forecast the extreme event indicator in the future $E(t+\delta t)$ after a time horizon $\delta t>0$ using $\boldsymbol{\hat{Q}}^{-}$ as the input to the model $\hat{f}$ such that
$\hat{E}(t+\delta t)=\hat{f}\left(\boldsymbol{\hat{Q}}^{-}\right),$ (3)
where $\hat{E}$ is the modeled extreme event indicator, which might differ from $E$. The performance of the model can be evaluated using the probability of error
$P_{e}(\boldsymbol{\hat{Q}}^{-},\hat{f})=\text{Probability}(\hat{E}\neq E)=P(\hat{f}(\boldsymbol{\hat{Q}}^{-})\neq E).$ (4)
@@ 2.1 Modeled extreme event indicator
Mispredicted extreme events can manifest as either false positives, $P(\hat{E}=1,E=0)$, or false negatives, $P(\hat{E}=0,E=1)$. However, these two types of errors can bear significantly different consequences. For instance, incorrectly predicting a hurricane (false positive) might be inconvenient but acceptable; however, failing to predict one (false negative) can be catastrophic. To accurately reflect the distinct impact of false positive and negative, we introduce the cost-sensitive probability of error:
@@
where $c^{+}>0$ and $c^{-}>0$ are the false positive and negative cost weighting factors, respectively. These factors reflect the relative severity of each type of error, and their values are selected according to the specific prediction task. In the case of extreme event prediction, the value of $c^{-}$ is often larger than $c^{+}$. This choice is driven by the understanding that non-extreme events occur more frequently than extreme ones. Prediction models with equal costs ($c^{+}=c^{-}$) are inclined to favor the majority class of non-extreme events. By imposing higher penalties on false negative errors ($c^{-}>c^{+}$), we steer the prediction model to focus more on accurately identifying the critical, but less frequent, extreme events. The value of the factors $c^{-}$ and $c^{+}$ is arbitrary, and only their relative magnitude matters. Consequently, $c^{-}$ and $c^{+}$ can be scaled in different manners. To guarantee that the model with the minimum probability of error yields $P_{e}^{c}<1/2$, we choose $21/c^{+}+1/c^{-}=$ 2 [see A for more details].
2.2 Minimum cost-sensitive probability of error
The goal is to estimate the minimum cost-sensitive probability of error given the observable $\boldsymbol{\hat{Q}}^{-}$ over all possible models $\hat{f}$,
@@ 2.2 Minimum cost-sensitive probability of error
The minimum cost-sensitive probability of error attainable by any model is [see proof in A]
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]=\sum_{\boldsymbol{\hat{q}}^{-}}I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),$ (7)
where $\boldsymbol{\hat{q}}^{-}$ is a particular state (i.e., value) for $\boldsymbol{\hat{Q}}^{-}$, $P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ is the probability of $\boldsymbol{\hat{Q}}^{-}$ taking the value $\boldsymbol{\hat{q}}^{-}$, and $I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ is the minimum probability of error for the state $\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}$:
where $P(E\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ is the probability of $E$ conditioned on $\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}$. The minimum error given by Eq. (7) is the consequence of the unavoidable uncertainty intrinsic to chaotic systems. This uncertainty arises from the lack of knowledge about the variables (e.g., unobserved variables and/or those observed for a limited amount of time) and errors in the initial condition values (e.g. finite precision), which transcend the predictive capabilities of any model.
Equation (7) provides the precise limit for extreme event forecasting; however, its application to the development, optimization, and evaluation of models for extreme event prediction is challenging due to its non-convex nature. This motivates the derivation of information-theoretic lower and upper bounds for $P_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-})$ that are more amenable in terms of applications and interpretation. For example, obtaining $P_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-})$ reliably from Equation (7) may not be possible in situations where, on the other hand, information-theoretic quantities can be efficiently calculated using estimators paninski2003 . Even when Equation (7) can be evaluated accurately, its manipulation becomes challenging in the context of model development due to the non-linearity introduced by the $\min(\cdot)$ operator hastie2009 . In such instances, using an information-theoretic formulation of the error facilitates the optimization of model parameters. Information theory can also be employed for feature selection, specifically identifying the input variables that most significantly aid in predicting extreme events guyon2003 . Additionally, the sources of error contributing to $P_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-})$ are more easily interpreted in terms of information rather than probabilities, since the former adheres to the properties of additivity and the chain rule cover2006 . In the next section, we derive lower and upper bounds for Eq. (7) using the framework of information theory.
2.3 Information-theoretic bounds for minimum probability of error
@@
The key idea to derive the information-theoretic bounds is that the prediction of extreme events can be intuitively understood as an information transmission process, where information from the current observable state is conveyed to predict the future state (LozanoDuran2022, ). If the forecast is treated as a noisy channel, then the Fano’s (Fano1961, ) and Hellman’s (hellman1970probability, ) inequalities provide the foundations for deriving lower and upper bounds on the minimum probability of error in the transmission of discrete messages. We measure the uncertainty in the extreme event indicator $E$ given the information from the observable $\boldsymbol{\hat{Q}}^{-}$ using the cost-sensitive, conditional Renyi entropy (renyi1961measures, ; Zhao2013, )
$H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ $=\sum_{\boldsymbol{\hat{q}}^{-}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),$ (9)
where
$h_{\alpha}^{c}(p)=h_{\alpha}\left(\min\{c^{-}p,c^{+}(1-p)\}\right)$
is the cost-sensitive binary Renyi entropy function of order $\alpha>0$, with
$h_{\alpha}(p)=\lim_{\gamma\rightarrow\alpha}\frac{1}{1-\gamma}\log_{2}\left(p^{\gamma}+(1-p)^{\gamma}\right).$
Equation (9) quantifies the additional information required to determine the outcome of $E$ given the information in $\boldsymbol{\hat{Q}}^{-}$ accounting for the weighting factors $c^{-}$ and $c^{+}$. It is useful to interpret $H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ as the uncertainty in $E$ after conducting the ‘measurement’ of $\boldsymbol{\hat{Q}}^{-}$. If $E$ and $\boldsymbol{\hat{Q}}^{-}$ are independent random variables, then $H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=H_{\alpha}^{c}(E)$, i.e., knowing $\boldsymbol{\hat{Q}}^{-}$ does not reduce the uncertainty in $E$. In this case, $\boldsymbol{\hat{Q}}^{-}$ is not a useful observable for forecasting $E$. Conversely, if knowing $\boldsymbol{\hat{Q}}^{-}$ provides the knowledge to completely determine $E$, then $H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=0$, i.e., there is no uncertainty in $E$ given $\boldsymbol{\hat{Q}}^{-}$, and the $\boldsymbol{\hat{Q}}^{-}$ can potentially predict $E$ with no error. The order $\alpha$ determines the extent to which different probabilities influence the uncertainty, with larger values of $\alpha$ giving greater weight to higher probabilities. For $c^{+}=c^{-}=\alpha=1$, $H_{1}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ is equal to the classic Shannon conditional entropy (Shannon1948, ), which is a concave function in the conditional distribution, making it well-suited for optimization tasks.
The minimum probability of error can be lower and upper bounded as a function of the cost-sensitive conditional Rényi entropy [see proof in A]
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-})\leq P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})\leq P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-}),$ (10)
where the lower and upper bounds are
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-})=h_{\alpha}^{-1}\left(H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)$ (11)
$P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-})=\min\left\{\frac{1}{2}H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),C\right\},$
and $C=\min\left\{c^{-}P(E=1),c^{+}(1-P(E=1))\right\}.$ Eq. (10) is valid for $0<\alpha\leq 2$, with the tightest bounds achieved for $\alpha=2$, i.e., the quadratic conditional entropy $H_{2}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$. Nonetheless, maintaining the more general formulation with $\alpha$ is beneficial, as it establishes a relationship between error and information within the context of different entropies. It is worth noting that the minimum error in Eq. (7) and the information-theoretic bounds in Eq. (10) hold for any value of the thresholding parameter $\eta$ defining the cutoff for extreme events. Furthermore, Eq. (7) and Eq. (10) are generally applicable to the prediction of any binary events, whether they are extreme or not.
A corollary from the conditional entropy inequality (a.k.a. information can’t hurt(cover2006, ) is that incorporating additional time lags into the vector of observables can decrease (but never increase) the minimum probability of error [see proof in B]
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-l})\leq P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-p}),\ \text{for}\ l>p,$ (12)
$P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-l})\leq P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-p}),\ \text{for}\ l>p,$
where $l$ and $p$ denote the number of time lags in $\boldsymbol{\hat{Q}}^{-l}$ and $\boldsymbol{\hat{Q}}^{-p}$, respectively, i.e., $\boldsymbol{\hat{Q}}^{-l}=[\boldsymbol{\hat{Q}}(t),\boldsymbol{\hat{Q}}(t-\delta t_{1}),\dots,\boldsymbol{\hat{Q}}(t-\delta t_{l})]\pm\delta\boldsymbol{\hat{Q}}^{-l}$ (and similarly for $\boldsymbol{\hat{Q}}^{-p}$). The inequality presented in Eq. (12) is particularly useful in scenarios where not every variable of the system is directly observable. In such instances, it is still possible to lower the minimum probability of error by employing additional time-lagged measurements of the observed variables. This result can be connected to Takens’ embedding theorem (Takens1981, ), whereby the dynamics of a dynamical system can be captured by embedding a sequence of past observations into a higher-dimensional space. The latter is consistent with the decrease in the probability of error from Eq. (12). Takens’ embedding theorem also states that the delayed-embedding phase space of $\boldsymbol{\hat{Q}}^{-l}$ is topologically equivalent to the original phase space of the full dynamical system ($\boldsymbol{Q}$) for a non-degenerate, noise-free observable with $l>2d_{A}$, where $d_{A}$ is the dimension of the attractor. From an information-theoretic viewpoint, this implies that $\boldsymbol{\hat{Q}}^{-l}$ provides the same information as the state vector of the full system $\boldsymbol{Q}$ when there is no uncertainty in the observable ($\delta\boldsymbol{\hat{Q}}^{-l}=\boldsymbol{0}$) and $l>2d_{A}$. Under those conditions, $P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-l})=P_{e,\text{min}}^{c}(\boldsymbol{Q})=0$ and exact predictions are possible.
The minimum probability of error from Eq. (7) and the information-theoretic bounds from Eq. (10) are derived assuming the discretization of the continuous variable $\boldsymbol{\hat{Q}}$. This reflects the inherent uncertainty in the initial conditions, which is captured by the term $\delta\boldsymbol{\hat{Q}}^{-}>\boldsymbol{0}$. Nonetheless, the same inequality holds when $\boldsymbol{\hat{Q}}$ is assumed to be a continuous variable by replacing the $H_{\alpha}^{c}$ with its continuous extension [see C]. In that case, there is no uncertainty in the value of $\boldsymbol{\hat{Q}}$, and we can take $\delta\boldsymbol{\hat{Q}}^{-}=\boldsymbol{0}$. It can also be shown that the right-hand side of Eq. (7) and the bounds in Eq. (11) converge to their continuous counterparts when the partition size used to discretize $\boldsymbol{\hat{Q}}$ is refined towards zero. Here, we focus on the discrete case, as in practical applications there is always some degree of uncertainty in $\boldsymbol{\hat{Q}}$.
2.4 Sources contributing to the probability of error